OpenStudy (anonymous):

if a 2gm hydrocarbon is oxidized,,we get 4.5 gm H2O and 5.5 gm CO2...what is the moleculer formula of the hydrocarbon????

OpenStudy (anonymous):

is there any body to help??????

OpenStudy (callisto):

Not sure if I am on the right path. Let \(C_xH_y\) be the hydrocarbon. The equation for reaction is \[C_xH_y + (x+\frac{y}{2}) O_2 \rightarrow \frac{y}{2}H_2O +x CO_2\] No. of mole of \(CO_2\) = x = \(\frac{5.5}{12.0+16.0\times 2}\) x =... No. of mole of \(H_2O\) = \(\frac{y}{2}=\frac{4.5}{1.0\times 2+16.0}\) y = ... Then find the LCM of x and y so that x and y are integers, and you'll get the molecular formula.